### v1

# NbBench: Benchmarking Language Models for Comprehensive Nanobody Tasks

Abstract
Nanobodies—single-domain antibody fragments derived from camelid heavy-chain-only antibodies—exhibit unique advantages such as compact size, high stability, and strong binding affinity, making them valuable tools in therapeutics and diagnostics. While recent advances in pretrained protein and antibody language models (PPLMs and PALMs) have greatly enhanced biomolecular understanding, nanobody-specific modeling remains underexplored and lacks a unified benchmark. To address this gap, we introduce NbBench, the first comprehensive benchmark suite for nanobody representation learning. Spanning eight biologically meaningful tasks across nine curated datasets, NbBench encompasses structure annotation, binding prediction, and developability assessment. We systematically evaluate eleven representative models—including general-purpose protein LMs, antibody-specific LMs, and nanobody-specific LMs—in a frozen setting. Our analysis reveals that antibody language models excel in antigen-related tasks, while performance on regression tasks such as thermostability and affinity remains challenging across all models. Notably, no single model consistently outperforms others across all tasks. By standardizing datasets, task definitions, and evaluation protocols, NbBench offers a reproducible foundation for assessing and advancing nanobody modeling. All datasets and source code used in our benchmark are publicly available at https://github.com/ZHymLumine/NbBench.
1 Introduction
Antibodies are essential components of the adaptive immune system, capable of recognizing and neutralizing diverse antigens such as viruses, bacteria, and toxins lu2020development . This high specificity has made antibodies indispensable in therapeutics, diagnostics, and molecular biology nelson2010development . Structurally, conventional antibodies are composed of two heavy and two light chains forming a Y-shaped molecule, with the antigen-binding regions located within the variable domains schroeder2010structure .
Nanobodies (also known as VHHs or single-domain antibodies) are the antigen-binding domains of heavy-chain-only antibodies found in camelids el2025nanobody . Unlike conventional antibodies, nanobodies consist of a single variable domain and lack light chains, yet retain high binding affinity and specificity muyldermans2013nanobodies ; jovvcevska2020therapeutic . Their simplified architecture confers several advantages: small size ( 15 kDa), high solubility, strong thermal and chemical stability, and the ability to access hidden epitopes el2025nanobody ; DeMeyer2014NanobodybasedPA ; van2016nanobodies . These features make nanobodies attractive for applications in molecular imaging, therapy, and synthetic biology alexander2024discovery ; bao2021nanobody .
Despite these advantages, nanobodies pose distinct modeling challenges. Their lack of paired light chains, longer and more flexible CDR3 loops muyldermans2013nanobodies , and smaller structural databases make it difficult to apply conventional antibody modeling techniques steeland2016nanobodies . While large datasets like the Observed Antibody Space (OAS)kovaltsuk2018observed ; olsen2022observed exist for conventional antibodies, nanobody-specific resources such as INDIdeszynski2022indi and VHHCorpus-2M tsuruta2024sars remain relatively limited.
Recent advances in protein language models (PLMs), such as ESM lin2023evolutionary , ProtBert elnaggar2021prottrans , and MSA Transformer rao2021msa , have demonstrated strong performance on structure and function prediction tasks rao2019evaluating ; madani2020progen ; meier2021language . These models leverage large-scale self-supervised training on protein sequences to learn contextual representations that capture structural and evolutionary information. Building on this, antibody-specific language models (PALMs) like AntiBERTy ruffolo2021deciphering , AbLang olsen2022ablang , AntiBERTa2 barton2024enhancing , and IgBert kenlay2024large have been developed to model immunoglobulin sequences with higher accuracy. However, their applicability to nanobody modeling remains largely unexplored, as they are predominantly trained and evaluated on conventional antibodies.
More recently, nanobody-specific models such as NanoBERT hadsund2024nanobert , NanoBERTa-ASP li2024nanoberta , DeepNano deng2024nanobody , and VHHBERT tsuruta2024sars have emerged. However, these models are typically developed for narrow tasks and use inconsistent datasets and evaluation metrics, making fair comparisons difficult. Furthermore, most prior work does not assess generalization to unseen nanobody types or functions.
To address these limitations, we introduce NbBench, the first comprehensive benchmark suite for nanobody modeling. NbBench spans eight biologically relevant tasks across nine curated datasets, covering structural annotation, sequence generation, function prediction, and developability assessment. It provides standardized task definitions, data splits, and evaluation metrics to facilitate fair, reproducible comparisons between general-purpose, antibody-specific, and nanobody-specific language models. Using NbBench, we evaluate eleven pretrained models and analyze their generalization performance across task types, revealing key insights into the capabilities and limitations of current language models for nanobody modeling.
Overall, our contributions can be summarized as follows:
  • We introduce NbBench, the first comprehensive benchmark for nanobody modeling, spanning eight biologically meaningful tasks across nine curated datasets.
  • We systematically evaluate eleven pretrained protein, antibody, and nanobody language models, revealing their generalization capabilities and failure modes across structure, function, and developability tasks.
  • We establish a standardized evaluation protocol for nanobody tasks, enabling fair model comparison and providing actionable insights for future model development and deployment.
Figure 1: Overview of NbBench. (a) Eight tasks spanning token- and sequence-level classification and regression. (b) Data sources include INDI, SAbDab, PLAbDab-nano, and task-specific datasets. (c) 10+ baselines from pretrained protein, antibody, and nanobody language models. (d) Visual representations of tasks grouped into structure understanding, binding, and biophysical and functional property prediction. (e) Standard evaluation metrics and clustering-based data splits.
2 Related Works
Pretrained Protein and Antibody Language Models (PPLMs and PALMs).Protein language models such as ESM-2 lin2023evolutionary , ProtBert elnaggar2021prottrans , and MSA Transformer rao2021msa have demonstrated strong transfer performance on structure and function prediction tasks rao2019evaluating ; madani2020progen ; meier2021language ; chen2023structure ; zhang2025hyena . These models leverage large-scale self-supervised learning and evolutionary context to build general-purpose protein representations. In parallel, antibody-specific language models (PALMs) have been introduced to model immunoglobulin sequences more effectively. AntiBERTy ruffolo2021deciphering was the first such model trained on 588 million heavy/light chain sequences from the OAS database, revealing that antibodies can be embedded along trajectories resembling affinity maturation. AntiBERTa leem2022deciphering was fine-tuned for paratope prediction, while AbLang olsen2022ablang trained separate models on heavy (AbLang-H) and light (AbLang-L) chains, outperforming ESM-1b in recovering masked residues. BERT-DS porebski2024rapid and AntiBERTa2 barton2024enhancing further expanded training datasets to tens or hundreds of millions of antibody sequences, with AntiBERTa2 incorporating multimodal contrastive learning between sequence and structure. IgBert kenlay2024large , initialized from ProtBert, was trained on over 2 million paired antibody sequences. EATLM wang2023pre introduced evolution-guided pretraining objectives, such as germline ancestor prediction and mutation site masking, and also proposed the ATUE benchmark. Despite this rapid progress, most PALMs have focused on conventional antibodies and are primarily evaluated on limited downstream tasks such as paratope prediction, sequence recovery, or general binding classification.
Nanobody modeling and current limitations.While antibody-specific language models have advanced rapidly, their applicability to nanobody modeling remains largely unexplored. Nanobodies—single-domain fragments derived from camelid antibodies—have emerged as promising tools for diagnostics, therapeutics, and molecular imaging salvador2019nanobody ; alexander2024discovery ; bao2021nanobody . Their compact size and ability to bind hidden epitopes make them attractive alternatives to conventional antibodies salvador2019nanobody ; liu2025unveiling . Classical in silico approaches—such as docking, molecular dynamics, and epitope mapping—have long been used to analyze nanobody-antigen interactions el2025nanobody ; joubbi2024antibody , but are often resource-intensive and lack scalability. Structure prediction tools like AlphaFold2 jumper2021highly and OmegaFold wu2022high improve modeling accuracy, yet still rely on structural inputs or evolutionary features valdes2023structural . Although a few deep learning models such as NanoBERT hadsund2024nanobert , NanoBERTa-ASP li2024nanoberta , and VHHBERT tsuruta2024sars have been proposed for thermostability, paratope, or binding prediction, each is designed for a narrow task and evaluated in isolation. A unified framework for evaluating nanobody modeling remains missing.
Benchmarking efforts and the need for NbBench.Large-scale benchmarks have played a critical role in evaluating protein and antibody language models. TAPE rao2019evaluating introduced five foundational tasks, such as secondary structure and stability prediction, to assess general-purpose protein representations. PEER xu2022peer expanded to multi-task evaluation across sequence- and token-level objectives, while FLIP dallago2021flip focused on mutational fitness landscapes to benchmark protein engineering. More recently, ATUE wang2023pre proposed the first antibody-specific benchmark, covering four real-world tasks—antigen binding prediction mason2021optimization , paratope prediction liberis2018parapred , B cell maturation analysis mroczek2014differences , and SARS-CoV-2 antibody discovery raybould2021cov ; zaslavsky2022disease —designed to assess the biological understanding and evolutionary awareness of pre-trained antibody models. Despite these advances, all existing benchmarks share several limitations in the context of nanobody modeling. First, they primarily focus on general proteins or conventional antibodies, with little attention to the structural and functional uniqueness of nanobodies. Second, tasks are often narrow in scope and evaluated in isolation, preventing consistent comparison across models. Third, none of the benchmarks provide a unified evaluation across the diverse range of nanobody-relevant tasks such as structure modeling, binding function, and developability. To address these gaps, we introduce NbBench, the first multi-task benchmark suite specifically designed for nanobody modeling. NbBench standardizes datasets, metrics, and splits across eight biologically meaningful tasks, enabling systematic and fair evaluation of both general-purpose and antibody-specific language models in nanobody-centric scenarios.
3 Benchmark Tasks
NbBench consists of 8 tasks across 10 datasets grouped into three categories: structure and region understanding, binding prediction, and biophysical property prediction. These tasks cover token-level and sequence-level classification, regression, and sequence generation with an overview provided in Table 1.
3.1 Structure and Region Understanding
Variable Region Classification (VRCls) classifies each residue in a nanobody sequence into structural regions: framework regions (FRs) or one of the three complementarity-determining regions (CDR1, CDR2, CDR3). This task is formulated as an amino acid-level classification problem, where the model must predict a categorical label for each amino acid in the sequence based on its structural and functional role. We collect data from the INDI database deszynski2022indi , using the abgenbank, manual, and structure subsets for training, and the patent subset for testing. We remove test sequences that are exact duplicates of those in the training set, and ensure that the average sequence similarity between training and test sets remains at 75%.
Evaluation metric: accuracy, precision, recall, and F1 score.
Biological impact: The three complementarity-determining regions (CDRs) are the primary determinants of antigen binding, while the surrounding framework regions (FRs) maintain the structural integrity required for proper presentation of the CDRs. However, the activity of a nanobody is not only determined by the CDR sequences themselves; the three-dimensional structure and interactions between all variable regions, including FRs and CDRs, play a crucial role in antigen recognition and binding nakanishi2008critical . Therefore, accurately classifying every residue in the variable region—whether it belongs to a framework or a CDR—is essential. This comprehensive annotation enables more effective downstream applications such as CDR grafting, humanization, affinity maturation, and structure-aware generative design, and improves the interpretability and reliability of nanobody models.
CDR Infilling (CDRInf) reconstructs masked CDR regions within a nanobody sequence. Given a sequence where the CDRs are masked or removed, the model is expected to generate plausible CDR sequences that complete the nanobody. This task is formulated as a token-level sequence generation problem. The dataset and train-test split are consistent with the CDR classification task. We use the same subsets from INDI and apply the sequence similarity-based splitting strategy.
Evaluation metric: Exact match accuracy (hard recovery) and similarity score based on BLOSUM62 (soft recovery).
Biological impact: CDR infilling evaluates the model’s ability to generate diverse yet functional antigen-binding regions within a fixed nanobody scaffold. Success on this task suggests a capacity to learn structure-function relationships that could be used in de novo nanobody design, especially in scaffold-constrained generative pipelines or CDR-optimization loops.
3.2 Binding Prediction
Antigen-Nanobody Binding Prediction (AgNbBind) predicts whether a nanobody can bind to a given antigen. This is formulated as a binary classification task. We include three separate datasets, each corresponding to a different antigen: (1) SARS-CoV-2 tsuruta2024sars and (2) human IL-6 (hIL6) tsuruta2023avida .
Evaluation metric: AUROC (area under the receiver operating characteristic), AUPRC (area under the precision-recall curve), accuracy, precision, recall and F1 score.
Biological impact: Predicting binding interactions is critical for therapeutic antibody discovery. It helps prioritize candidates with high affinity and broad neutralization potential, and accelerates lead selection and optimization.
Paratope Prediction (Paratope) identifies which residues in the nanobody are directly involved in antigen binding. This is framed as a sequence labeling task, where the model predicts a binary label for each residue in the nanobody, indicating whether it is part of the paratope.
Evaluation metric: AUROC, AUPRC, accuracy, precision, recall and F1 score.
Biological impact: Understanding which residues form the paratope reveals how nanobodies interact with antigens. This insight supports rational nanobody engineering, such as improving binding affinity through site-directed mutagenesis or redesigning specific binding loops.
3.3 Biophysical and Functional Property Prediction
Thermostability Prediction (Thermo) estimates the melting temperature (Tm) of nanobody sequences, which reflects their thermal stability. This task is framed as a regression problem, where the model is trained to predict a continuous Tm value from the nanobody sequence alone. We collect thermostability data from multiple curated sources, including NbThermo valdes2023nbthermo , TEMPRO alvarez2024tempro , and NanoMelt ramon2025prediction , which report experimentally measured or computationally inferred Tm values for diverse nanobodies. We apply sequence similarity-based splitting to ensure low redundancy between training and test sets, and to simulate generalization to novel scaffolds.
Evaluation metric: Spearman correlation, $R^{2}$, RMSE (root mean square error), and MAE (mean absolute error).
Biological impact: Thermostability is a key developability attribute for therapeutic and diagnostic nanobodies. Higher Tm values are associated with better solubility, longer shelf life, and resilience in extreme environments, making thermostability prediction an important task for early-stage candidate screening and protein engineering.
Polyreactivity Prediction (PolyRx) predicts whether a nanobody exhibits polyreactivity, defined as nonspecific binding to off-target proteins or biomolecules. This task is formulated as a binary classification problem to distinguish between polyreactive and non-polyreactive sequences. We collect data from two prior studies harvey2022silico ; zhou2024gp , which experimentally measured and annotated nanobody polyreactivity using a naïve synthetic camelid antibody fragment library.
Evaluation metric: AUROC, AUPRC, accuracy, precision, recall and F1 score.
Biological impact: Polyreactivity is a major liability in therapeutic antibody development, as it can cause off-target effects, reduce specificity, and lead to misleading experimental results. Predicting and minimizing polyreactivity during early-stage screening improves the reliability and clinical developability of nanobody-based drugs.
[…]
Evaluation metric: accuracy, precision, recall, and F1 score.
Biological impact:Correctly identifying antibody types is essential for immune repertoire profiling, especially in mixed or synthetic libraries. Differentiating nanobodies (e.g., VHH and VNAR) from conventional antibodies enables more targeted therapeutic and diagnostic engineering.
VHH Affinity Prediction (Affinity) estimates the antigen-binding strength of a VHH nanobody based on its amino acid sequence. This is formulated as a regression task, where the output is a continuous affinity-related score derived from high-throughput biopanning experiments. We collect data from the study by tuvcs2024extensive , which applied deep sequencing to phage-displayed nanobody libraries across multiple rounds of selection. Binding scores are computed based on the relative enrichment of sequences between non-specifically bound phages (N.S. phages) and eluted phages in the 4th round of selection. Specifically, the score for each sequence is defined as:
[…]
Evaluation metric: Spearman correlation, $R^{2}$, RMSE and MAE.
Biological impact: Accurate prediction of nanobody affinity enables virtual screening of large candidate pools, reducing the need for costly wet-lab experiments. This task supports sequence design strategies such as multi-objective optimization.
3.4 Data Splitting Strategy
Dataset Splitting via Sequence Clustering.
To ensure a fair and biologically meaningful evaluation, we adopt a clustering-based strategy for splitting datasets using MMseqs2 steinegger2017mmseqs2 with a 70% sequence identity threshold, except for the Antigen-Nanobody Binding Prediction task.
We first cluster the entire sequence database such that highly similar sequences (with identity $\geq 70\%$) are grouped together. To encourage diversity, 20% of the sequences are sampled from low-frequency clusters to form the test set. The remaining sequences are split into 70% for training and 10% for validation. For regression tasks such as Thermo and Affinity, we further applied stratified sampling based on binned target values to ensure that the label distribution is consistent across training, validation, and test sets.
To verify the effectiveness of the split, we compute the average pairwise identity between the training and test sets. This ensures that the test set contains sequences that are dissimilar to those seen during training, simulating realistic generalization scenarios and minimizing potential data leakage.
This approach reduces redundancy across splits and prevents data leakage. More importantly, it simulates realistic generalization scenarios in which models encounter novel, non-homologous sequences. This is particularly critical for tasks such as nanobody-antigen interaction prediction, where high sequence similarity may lead to inflated evaluation metrics if not properly controlled.
Antigen-Nanobody Binding Split.
For the Antigen-Nanobody Binding Prediction task, we use a task-specific split based on known biological variations in the antigens: SARS-CoV-2: The training set includes wild-type (WT), mild variants (e.g., D614G), conserved regions (S2 domain), and immune escape mutations (PMS). The validation set contains unseen variants from the Kappa and Lambda lineages for hyperparameter tuning. The test set consists of major variants of concern (Alpha, Beta, Delta, Omicron), simulating real-world generalization to novel strains. hIL6: The training set uses the wild-type antigen; the validation set consists of 5 randomly selected mutants, and the test set includes the remaining 25 mutants.
Table 1: Overview of benchmark tasks for nanobody modeling. Token-level tasks require a label for each amino acid; Sequence-level tasks assign a single label per nanobody sequence. Cls and Reg denote classification and regression tasks, respectively. Task suffixes: -seq indicates sequence clustering-based splits; -tm and -score denote stratified splits based on melting temperature and binding affinity scores, respectively.
Task #Train/Val/Test Type Granularity Metrics Source Structure and Region Understanding VRCls 13,703/3,426/2,847 Multi-class Cls Token Accuracy, Precision, Recall, F1 INDI deszynski2022indi CDRInf 13,703/3,426/2,847 Multi-class Cls Token Exact Match, BLOSUM62 Similarity INDI deszynski2022indi Binding Prediction SARS-CoV-2 27,606/6,706/38,958 Binary Cls Sequence AUROC, AUPRC, Accuracy, Precision, Recall, F1 AVIDa-SARS-CoV-2 tsuruta2024sars hIL6 13,020/11,637/449,234 Binary Cls Sequence AUROC, AUPRC, Accuracy, Precision, Recall, F1 AVIDa-hIL6 tsuruta2023avida Paratope 851/139/240 Binary Cls Token AUROC, AUPRC, Accuracy, Precision, Recall, F1 SAbDab wilton2018sdab Biophysical and Functional Property Prediction Thermo-seq 522/92/147 Regression Sequence Spearman, $R^{2}$, RMSE, MAE NbThermo/NanoMelt valdes2023nbthermo ; ramon2025prediction Thermo-tm 396/57/114 Regression Sequence Spearman, $R^{2}$, RMSE, MAE NbThermo/NanoMelt ramon2025prediction PolyRx 101,854/14,613/25,007 Binary Cls Sequence AUROC, AUPRC, Accuracy, Precision, Recall, F1 Harvey/GP-nano harvey2022silico ; zhou2024gp NbType 12,647/1,960/3,557 Multi-class Cls Sequence Accuracy, Precision, Recall, F1 PLAbDab-nano/AbNativ gordon2025plabdab ; ramon2024assessing Affinity-seq 8,888/1,302/2,547 Regression Sequence Spearman, $R^{2}$, RMSE, MAE MOQA tuvcs2024extensive Affinity-score 8,915/1,274/2,548 Regression Sequence Spearman, $R^{2}$, RMSE, MAE MOQA tuvcs2024extensive
4 Benchmark Models
To comprehensively evaluate nanobody modeling performance, we select a suite of baseline language models covering general protein, antibody, and nanobody-specific pre-training domains. These models vary in architecture, pre-training corpora, and design philosophies. As a practical constraint, we only include models with publicly available weights to ensure reproducibility. Unless otherwise noted, all models are evaluated in a frozen setting using amino acid-level tokenization.We describe each model in detail below and present a summary in Table 2.
Protein Language Models. (1) ProtBert elnaggar2021prottrans is a BERT-based model pre-trained on 216 million sequences from UniRef100. It serves as a general-purpose protein language model. (2–3) ESM-2 lin2023evolutionary is a Transformer-based model trained on 65 million sequences from UniRef50. We evaluate two variants: ESM-2 150M and ESM-2 650M. These models help assess the transferability of general protein representations to nanobody tasks.
Antibody Language Models. (4) AbLang-H olsen2022ablang is a RoBERTa-based model trained on 14 million heavy-chain sequences from the OAS database. (5) AntiBERTy ruffolo2021deciphering is a BERT-based model pre-trained on 588 million antibody sequences from OAS. (6) AntiBERTa2 barton2024enhancing is a RoFormer-based model trained on 824 million sequences from OAS and a proprietary dataset. (7) AntiBERTa2-CSSP barton2024enhancing is a multimodal variant of AntiBERTa2 that incorporates antibody structure via contrastive sequence–structure pre-training. (8) IgBert kenlay2024large is initialized from ProtBert and further trained on 2 billion unpaired heavy&light chain sequences and 2 million paired antibody sequences.
Nanobody Language Models. (9) NanoBERT hadsund2024nanobert is a RoBERTa-based model trained specifically on nanobody sequences to explore mutational space and thermostability. (10) VHHBERT tsuruta2024sars is a RoBERTa-based model pre-trained on 2 million camelid nanobody sequences from VHHCorpus-2M. It uses extended positional embeddings to handle long CDR3 loops common in nanobodies.
To ensure fair comparison and isolate the representational quality of pretrained models, all models are evaluated in a frozen setting—i.e., their parameters remain unchanged—and only the final task-specific classifier (e.g., MLP) is trained during fine-tuning. This protocol allows us to assess how well each pretrained representation transfers to downstream nanobody tasks.
Table 2: Characteristics of pre-trained protein, antibody, and nanobody language models. “M” stands for million.
Model Number of Parameters (M) Pre-training Dataset Data Type Protein Language Models ProtBert elnaggar2021prottrans 420M UniRef100 suzek2007uniref (216M) General Protein ESM-2 (150M) lin2023evolutionary 150M UniRef50 suzek2007uniref (65M) General Protein ESM-2 (650M) lin2023evolutionary 650M UniRef50 suzek2007uniref (65M) General Protein Antibody Language Models AbLang-H olsen2022ablang 86M OAS kovaltsuk2018observed (14M heavy) Heavy Chain AbLang-L olsen2022ablang 86M OAS kovaltsuk2018observed (0.24M light) Light Chain AntiBERTy ruffolo2021deciphering 26M OAS kovaltsuk2018observed (588M) Heavy, Light Antibody AntiBERTa2 barton2024enhancing 203M OAS kovaltsuk2018observed + proprietary (824M) Heavy, Light Chain AntiBERTa2-CSSP barton2024enhancing 202M OAS kovaltsuk2018observed + structure contrastive (824M) Heavy, Light Chain IgBert kenlay2024large 420M OAS kovaltsuk2018observed + 2B unpaired + 2M paired Heavy, Light Chain Nanobody Language Models NanoBERT hadsund2024nanobert 15M INDI deszynski2022indi (10M) Heavy Nanobody VHHBERT tsuruta2024sars 86M VHHCorpus tsuruta2024sars (2M) Heavy Nanobody
5 Results
5.1 Task Pipeline
NbBench tasks fall into two main categories: sequence-level and token-level prediction. Each category includes two variants, depending on whether the task involves only nanobody sequences or both nanobody and antigen sequences. Antigen embeddings are consistently extracted using ESM-2 (650M), while nanobody sequences are encoded using each respective pretrained model under evaluation.
Sequence-Level Prediction. We consider two types of sequence-level tasks:
Nanobody-only Tasks: The model takes the nanobody sequence as input and uses the [CLS] token representation to generate a fixed-length embedding. This embedding is passed through a multilayer perceptron (MLP) for sequence-level classification or regression. Nanobody-Antigen Tasks: For tasks such as antigen-nanobody binding prediction, we first extract a global embedding of the antigen using ESM-2 (650M). This embedding is then concatenated with the [CLS] token from the nanobody sequence. The concatenated representation is pooled and processed through an MLP to generate the final prediction.
Token-Level Prediction. Token-level tasks also fall into two types:
Nanobody-only Tasks: The model produces token-level embeddings for each amino acid in the nanobody sequence. These embeddings are independently passed through an MLP to make token-wise predictions (e.g., for CDR classification and CDR infilling).
Nanobody-Antigen Tasks: In paratope prediction, we compute antigen embeddings using ESM-2 (650M), then concatenate the antigen representation with each position in the nanobody sequence. The resulting context-aware embeddings are mean pooled across antigen tokens and passed through multiple residual blocks to produce per-residue predictions on the nanobody.
5.2 Benchmark Results
We evaluated eleven frozen backbone models—including general protein language models, antibody models, and nanobody-specific models—on the twelve tasks in NbBench. The main evaluation metric for each task is shown in Table 3, while additional metrics such as precision, recall, and RMSE are provided in Appendix A.5. All experiments were run using three different random seeds, and we report the average performance with standard deviations.
Easy structure–type tasks. Variableregion classification (VRCls) and nanobody type classification (NbType) are essentially solved: every model exceeds 98% accuracy. Primary–sequence motifs already suffice to separate frameworks from CDRs and to tell VHH from VNAR, so these two tasks are no longer discriminative.
Antigen–related tasks favour antibody LMs. On SARS-2 and hIL6 binding as well as Paratope prediction, antibody LMs dominate; AntiBERTa2-CSSP reaches 0.92 AUROC for hIL6 and 0.94 for paratope. The large immunoglobulin corpora used during pretraining clearly encode crossmolecule interaction cues that general protein LMs—and the smaller nanobody LMs—do not capture.
Regression remains hard. For numerical traits such as Thermo stability and Affinity, all models perform modestly (best Spearman $\rho\approx 0.59$ for stability and <0.20<0.20 for affinity). Pure sequence embeddings—kept frozen here—appear insufficient to recover finegrained physicochemical signals without structural or experimental features.
No universal winner. Antibody LMs give the best overall balance for classification and binding; protein LMs lead on PolyRx. These results confirm that model performance is highly taskspecific and that future work should combine domainmatched pretraining with structureaware or parameterefficient fine tuning to push the remaining difficult fronts.
Table 3: Performance of various protein, antibody, and nanobody language models on the NbBench benchmark. Each value corresponds to the task’s main evaluation metric (e.g., Accuracy, AUROC, Spearman). Shading indicates top-3 performance for each task: darker is better. BR stands for BLOSUM62 Recovery, a soft recovery metric computed using BLOSUM62 substitution matrix. Mean (std) is reported for each experiment using three different seeds.
[…]
6 Conclusions and Limitations
In this work, we introduce NbBench, the first comprehensive benchmark suite for nanobody modeling. It spans eight biologically meaningful tasks—grouped into structure understanding, antigen binding, and biophysical property prediction—and covers ten rigorously curated datasets. To ensure fair and interpretable evaluation, we systematically assess eleven pretrained language models from the protein, antibody, and nanobody domains.
All models are evaluated in a frozen setting, where pretrained parameters remain fixed and only a lightweight classifier is trained on top. This design isolates the representational power of each model and avoids confounding factors from full fine-tuning, allowing us to fairly compare performance across diverse architectures and pretraining domains.
Our results highlight clear performance differences across model types. Antibody-specific models generally outperform others on antigen-related tasks, while nanobody-specific models show strength in thermostability and binding-related predictions. However, no single model dominates across all tasks, revealing that nanobody modeling remains a multifaceted challenge requiring tailored approaches.
NbBench provides a standardized and reproducible foundation for benchmarking nanobody models, but several limitations remain. First, we do not include large-scale LLMs such as GPT-4 or protein-specific foundation models requiring extensive resources, due to accessibility and reproducibility concerns. Second, the benchmark currently focuses on sequence-based tasks, given the limited availability of high-quality nanobody structural data. As more structures become available, future versions of NbBench could incorporate 3D structure-aware tasks such as docking, folding accuracy, or epitope prediction.
We hope NbBench catalyzes progress in nanobody representation learning, design, and downstream applications, and serves as a practical resource for both model development and rigorous evaluation in this growing domain.
Appendix A Appendix
A.1 Experimental Settings for Tasks
All tasks in NbBench are trained using a shared set of hyperparameters and training settings, as summarized in Table 4. All experiments were performed on a single NVIDIA A100 GPU with 80GB of memory.
Table 4: Training configuration for most of the benchmark tasks.
Config Value
optimizer AdamW
optimizer epsilon 1e-8
optimizer momentum $\beta_{1}=0.9$, $\beta_{2}=0.999$
weight decay 0.01
learning rate schedule linear decay
learning rate range [1e-5, 5e-3]
warmup steps 50
epochs 50
batch size 32
gradient accumulation steps 2
dtype float32
A.2 Data Preprocessing
We curated and preprocessed datasets for each nanobody-related task in NbBench to ensure reliable evaluation and reduce information leakage between training and testing splits.
Sequence Deduplication and Clustering.
To avoid data leakage, we removed 100% identical sequences between training and testing sets. We then applied MMseqs2 clustering with a minimum sequence identity threshold of 0.7 to split sequences into train, validation, and test sets. This ensured that sequences within the same cluster remained in the same split, minimizing cross-split similarity.
Variable Regions Classification and CDRs Infilling.
Training data were sourced from AbGenBank, manual curation, and structural databases, while testing sequences were extracted from patents. Each residue in the nanobody sequence was labeled according to its region (FR: 0, CDR1: 1, CDR2: 2, CDR3: 3), enabling token-level classification. For the CDRs infilling task, residues in the CDR1, CDR2, and CDR3 regions were masked by replacing them with hyphens (’-’), and language models were tasked with inferring the masked amino acids at these positions. This setup evaluates the models’ ability to reconstruct biologically meaningful variable regions based on the surrounding framework sequences.
Antigen-Nanobody Binding.
We collected binding pairs for SARS-CoV-2 and hIL6. For SARS-CoV-2, the training set includes wild type and variants with mild mutations, while the test set comprises Alpha, Beta, Delta, and Omicron variants to assess generalization. His-tags were removed from antigens such as hIL6 and hTNFa to ensure clean sequence input. Antigen embeddings were precomputed using ESM-2 (650M).
Epitope and Paratope Prediction.
Structural data from SAbDab-Nano were filtered to include only complexes with resolution < 3.0 Å. Residue pairs within 5 Å were labeled as interacting. Token-level binary classification was applied to the nanobody (paratope) sequences.
Stability and Polyreactivity.
For nanobody thermostability prediction, we compiled 764 sequences from multiple databases and stratified them by melting temperature (TM). The average sequence similarity between train and test sets was around 75%. For polyreactivity, high and low reactivity classes were separately clustered and merged afterward to ensure balanced representation and limited inter-class similarity.
Nanobody Type and Affinity prediction.
We classified antibodies into five types: VHH, VH, VNAR, Vkappa, and Vlambda, using balanced splits with inter-set similarity controlled (e.g., 76–80%). For affinity prediction, only sequences with binding scores $\geq 5$ or $\leq 1$ were retained for binary classification.
Similarity Control.
Across all tasks, MMseqs2’s "easy-search" module was used to compute the average similarity between training and testing sequences. All tasks reported train-test similarities between 73% and 84%, ensuring non-trivial generalization challenges.
A.3 Details for Antigen Binding and Paratope Prediction Tasks
In the antigen binding prediction task, the model input consists of a nanobody sequence and a precomputed antigen embedding. The nanobody sequence is encoded using a pretrained nanobody model, and the [CLS] token embedding is extracted from the final layer. Separately, the antigen embedding is generated using ESM-2 (650M) and mean-pooled along the sequence dimension. The pooled antigen embedding is then concatenated with the nanobody [CLS] embedding. The combined representation is passed through a linear classifier to predict binding outcomes, supporting both binary and multi-class classification, as well as regression settings depending on the task configuration.
For paratope prediction, the model receives the nanobody sequence along with an antigen embedding. The nanobody sequence is processed to produce token-level embeddings for each amino acid. Simultaneously, the antigen embedding is mean-pooled and expanded across the length of the nanobody sequence. At each amino acid position, the token embedding and the pooled antigen embedding are concatenated to form context-aware representations. These representations are further processed through three residual blocks and finally passed through a token-wise classifier to predict paratope labels for each residue.
A.4 BLOSUM62 Recovery
In the CDR Infilling task, we additionally report BLOSUM62 Recovery (BR), which assigns partial credit for biochemically similar substitutions, complementing the strict evaluation of Exact Match (EM).
Motivation and biological significance: The BLOSUM62 matrix reflects empirical substitution probabilities between amino acids observed in evolutionarily related proteins. Many amino acid substitutions are functionally tolerated because they preserve key chemical properties such as charge, hydrophobicity, or size. Therefore, even if a model’s prediction does not exactly match the ground-truth residue, it may still maintain biological function if the substitution is favorable according to BLOSUM62. BR captures this notion of functional robustness, providing a more biologically meaningful evaluation of sequence recovery, especially for applications like nanobody design, where partial recovery can still yield viable binding or stability.
Calculation: Given the ground-truth sequence $\mathbf{y}$ and the predicted sequence $\hat{\mathbf{y}}$, the BLOSUM62 score for each position is retrieved using the substitution matrix. The overall BLOSUM62 Recovery score is computed as the average substitution score across all valid positions:
$\text{BR}=\frac{1}{N}\sum_{i=1}^{N}\text{BLOSUM62}(y_{i},\hat{y}_{i})$
where $N$ denotes the number of valid amino acid positions (excluding padding tokens). In practice, scores are normalized if needed for comparability.
Furthermore, we separately compute token-level BR across the entire sequence and CDR-specific BR scores (for CDR1, CDR2, and CDR3), allowing finer-grained evaluation.
The pseudocode used to calculate BR is summarized below, and the full BLOSUM62 substitution matrix used for scoring is provided in Table 5.
  • For each sample, identify valid amino acid positions.
  • Partition the positions into CDR1, CDR2, and CDR3.
  • For each position, retrieve the BLOSUM62 score between the ground-truth and predicted residues.
  • Accumulate scores separately for the full sequence and each CDR region.
  • Report the mean BLOSUM62 score overall and per CDR.
Comparison with Exact Match (EM): While EM measures the fraction of residues that are predicted perfectly, it does not distinguish between near-correct and completely incorrect predictions. In contrast, BR provides a graded measure that rewards chemically plausible substitutions, thus offering a more nuanced assessment of model outputs.
Table 5: Complete BLOSUM62 substitution matrix.
A R N D C Q E G H I L K M F P S T W Y V B Z X * A 4 -1 -2 -2 0 -1 -1 0 -2 -1 -1 -1 -1 -2 -1 1 0 -3 -2 0 -2 -1 0 -4 R -1 5 0 -2 -3 1 0 -2 0 -3 -2 2 -1 -3 -2 -1 -1 -3 -2 -3 -1 0 -1 -4 N -2 0 6 1 -3 0 0 0 1 -3 -3 0 -2 -3 -2 1 0 -4 -2 -3 3 0 -1 -4 D -2 -2 1 6 -3 0 2 -1 -1 -3 -4 -1 -3 -3 -1 0 -1 -4 -3 -3 4 1 -1 -4 C 0 -3 -3 -3 9 -3 -4 -3 -3 -1 -1 -3 -1 -2 -3 -1 -1 -2 -2 -1 -3 -3 -2 -4 Q -1 1 0 0 -3 5 2 -2 0 -3 -2 1 0 -3 -1 0 -1 -2 -1 -2 0 3 -1 -4 E -1 0 0 2 -4 2 5 -2 0 -3 -3 1 -2 -3 -1 0 -1 -3 -2 -2 1 4 -1 -4 G 0 -2 0 -1 -3 -2 -2 6 -2 -4 -4 -2 -3 -3 -2 0 -2 -2 -3 -3 -1 -2 -1 -4 H -2 0 1 -1 -3 0 0 -2 8 -3 -3 -1 -2 -1 -2 -1 -2 -2 2 -3 0 0 -1 -4 I -1 -3 -3 -3 -1 -3 -3 -4 -3 4 2 -3 1 0 -3 -2 -1 -3 -1 3 -3 -3 -1 -4 L -1 -2 -3 -4 -1 -2 -3 -4 -3 2 4 -2 2 0 -3 -2 -1 -2 -1 1 -4 -3 -1 -4 K -1 2 0 -1 -3 1 1 -2 -1 -3 -2 5 -1 -3 -1 0 -1 -3 -2 -2 0 1 -1 -4 M -1 -1 -2 -3 -1 0 -2 -3 -2 1 2 -1 5 0 -2 -1 -1 -1 -1 1 -3 -1 -1 -4 F -2 -3 -3 -3 -2 -3 -3 -3 -1 0 0 -3 0 6 -4 -2 -2 1 3 -1 -3 -3 -1 -4 P -1 -2 -2 -1 -3 -1 -1 -2 -2 -3 -3 -1 -2 -4 7 -1 -1 -4 -3 -2 -2 -1 -2 -4 S 1 -1 1 0 -1 0 0 0 -1 -2 -2 0 -1 -2 -1 4 1 -3 -2 -2 0 0 0 -4 T 0 -1 0 -1 -1 -1 -1 -2 -2 -1 -1 -1 -1 -2 -1 1 5 -2 -2 0 -1 0 -4 -4 W -3 -3 -4 -4 -2 -2 -3 -2 -2 -3 -2 -3 -1 1 -4 -3 -2 11 2 -3 -4 -3 -2 -4 Y -2 -2 -2 -3 -2 -1 -2 -3 2 -1 -1 -2 -1 3 -3 -2 -2 2 7 -1 -3 -2 -1 -4 V 0 -3 -3 -3 -1 -2 -2 -3 -3 3 1 -2 1 -1 -2 -2 0 -3 -1 4 -3 -2 -1 -4 B -2 -1 3 4 -3 0 1 -1 0 -3 -4 0 -3 -3 -2 -2 0 -4 -3 -3 4 1 -1 -4 Z -1 0 0 1 -3 3 4 -2 0 -3 -3 1 -1 -3 -1 0 -1 -3 -2 -2 1 4 -1 -4 X 0 -1 -1 -1 -2 -1 -1 -1 -1 -1 -1 -1 -1 -1 -2 0 0 -2 -1 -1 -1 -1 -1 -4 -4 -4 -4 -4 -4 -4 -4 -4 -4 -4 -4 -4 -4 -4 -4 -4 -4 -4 -4 -4 -4 -4 -4 1
A.5 Full Results
In this section, we report the full evaluation results of each task in NbBench, grouped by task category. For each task, we provide a detailed performance table covering all evaluated models.
Structure and Region Understanding. We evaluate two tasks in this category: Variable Region Classification (VRCls) and CDR Infilling (CDRInf). Table 6 summarizes the performance on VRCls using standard classification metrics such as accuracy, precision, recall, and F1 score. For the CDRInf task, we evaluate both exact match (EM) and BLOSUM62 Recovery (BR) scores for each of the three CDRs and their aggregate, as shown in Table 7. These results allow us to assess both strict correctness and biochemical plausibility in recovered sequences.
Binding Prediction. We report model performance on three classification tasks: binding to SARS-CoV-2, binding to human IL-6 (hIL6), and paratope prediction. Each task is evaluated using AUROC, AUPRC, accuracy, precision, recall, and F1 score. Table 9 provides a comprehensive overview of results across these tasks. These metrics help evaluate how well models capture structural determinants of binding and functional interfaces.
Thermostability and Affinity Prediction. We assess model performance on four regression tasks: Thermo-seq and Thermo-tm (predicting thermal stability), Affinity-seq and Affinity-score (predicting binding affinity). Each task is evaluated using the Spearman correlation, $R^{2}$, RMSE, and MAE. Results are shown in the second half of Table 9, which has been reused for consistency. These evaluations capture how well models generalize to scalar-valued biophysical properties.
Polyreactivity Prediction. This classification task evaluates whether a nanobody exhibits nonspecific binding across diverse targets. As shown in Table 10, we report standard classification metrics (AUROC, AUPRC, accuracy, precision, recall, F1). Polyreactivity prediction is essential for screening nanobody candidates for developability.
Nanobody Type Prediction. We evaluate model accuracy in classifying the nanobody’s type (e.g., VHH from alpaca vs. VNAR from shark). Results in Table 11 show near-ceiling performance for several models, with high precision and recall across the board.
Table 6: Performance on the VRCls task (normalized to [0, 1]).
Model Accuracy Precision Recall F1 Score
ProtBert 0.9983 (0.0004) 0.9974 (0.0001) 0.9975 (0.0002) 0.9975 (0.0002)
ESM-2 (150M) 0.9989 (0.0004) 0.9982 (0.0004) 0.9987 (0.0005) 0.9985 (0.0001)
ESM-2 (650M) 0.9984 (0.0003) 0.9979 (0.0001) 0.9990 (0.0002) 0.9984 (0.0003)
AbLang-H 0.9981 (0.0001) 0.9964 (0.0002) 0.9959 (0.0004) 0.9962 (0.0003)
AbLang-L 0.9984 (0.0002) 0.9941 (0.0003) 0.9915 (0.0004) 0.9928 (0.0001)
AntiBERTy 0.9983 (0.0004) 0.9967 (0.0004) 0.9965 (0.0002) 0.9966 (0.0002)
AntiBERTa2 0.9989 (0.0005) 0.9986 (0.0002) 0.9972 (0.0001) 0.9979 (0.0001)
AntiBERTa2-CSSP 0.9983 (0.0004) 0.9968 (0.0003) 0.9977 (0.0004) 0.9972 (0.0004)
IgBert 0.9981 (0.0003) 0.9980 (0.0005) 0.9949 (0.0003) 0.9965 (0.0003)
NanoBERT 0.9947 (0.0004) 0.9916 (0.0003) 0.9885 (0.0004) 0.9900 (0.0003)
VHHBERT 0.9904 (0.0004) 0.9836 (0.0001) 0.9776 (0.0002) 0.9806 (0.0002)
Table 7: Performance on the CDRInf task. EM and BR stand for exact match and BLOSUM62 Recovery, respectively.
Model CDR1 EM CDR2 EM CDR3 EM CDRs EM CDR1 BR CDR2 BR CDR3 BR CDRs BR ProtBert 0.543 (0.0008) 0.533 (0.0002) 0.271 (0.0005) 0.406 (0.0004) 2.597 (0.0056) 2.403 (0.0021) 0.537 (0.0068) 1.525 (0.0039) ESM-2 (150M) 0.543 (0.0005) 0.531 (0.0006) 0.265 (0.0005) 0.402 (0.0003) 2.576 (0.0049) 2.400 (0.0024) 0.499 (0.0055) 1.500 (0.0034) ESM-2 (650M) 0.553 (0.0005) 0.539 (0.0003) 0.269 (0.0003) 0.408 (0.0002) 2.651 (0.0010) 2.464 (0.0021) 0.536 (0.0054) 1.553 (0.0029) AbLang-H 0.533 (0.0038) 0.524 (0.0013) 0.263 (0.0018) 0.397 (0.0020) 2.552 (0.0117) 2.340 (0.0108) 0.442 (0.0482) 1.452 (0.0290) AbLang-L 0.520 (0.0009) 0.516 (0.0006) 0.249 (0.0025) 0.385 (0.0011) 2.441 (0.0258) 2.292 (0.0073) 0.336 (0.0379) 1.360 (0.0144) AntiBERTy 0.541 (0.0014) 0.529 (0.0013) 0.266 (0.0020) 0.401 (0.0013) 2.560 (0.0088) 2.382 (0.0077) 0.514 (0.0335) 1.499 (0.0184) AntiBERTa2 0.538 (0.0006) 0.531 (0.0002) 0.263 (0.0006) 0.399 (0.0003) 2.539 (0.0043) 2.405 (0.0010) 0.494 (0.0021) 1.489 (0.0020) AntiBERTa2-CSSP 0.535 (0.0010) 0.528 (0.0015) 0.263 (0.0008) 0.398 (0.0010) 2.515 (0.0059) 2.385 (0.0118) 0.506 (0.0041) 1.484 (0.0041) IgBert 0.516 (0.0491) 0.502 (0.0526) 0.249 (0.0268) 0.380 (0.0389) 2.383 (0.3623) 2.190 (0.3770) 0.321 (0.2880) 1.311 (0.3290) NanoBERT 0.546 (0.0016) 0.530 (0.0030) 0.266 (0.0015) 0.403 (0.0011) 2.637 (0.0078) 2.402 (0.0210) 0.515 (0.0182) 1.524 (0.0115) VHHBERT 0.464 (0.0037) 0.490 (0.0014) 0.256 (0.0005) 0.367 (0.0007) 1.983 (0.0394) 2.101 (0.0104) 0.436 (0.0326) 1.244 (0.0219)
Table 8: Performance on SARS-CoV-2, hIL6 binding prediction, and paratope prediction tasks. Each task reports AUROC, AUPRC, Accuracy (Acc), Precision (Prec), Recall (Rec), and F1. The mean and standard deviation are reported for three different seeds.
Model SARS-2 Binding hIL6 Binding Paratope Prediction AUROC AUPRC Accu Prec Rec F1 AUROC AUPRC Acc Prec Rec F1 AUROC AUPRC Acc Prec Rec F1 ProtBert 0.852 (0.004) 0.785 (0.008) 0.847 (0.005) 0.845 (0.031) 0.565 (0.020) 0.676 (0.011) 0.878 (0.011) 0.578 (0.031) 0.973 (0.002) 0.888 (0.052) 0.373 (0.076) 0.519 (0.064) 0.916 (0.004) 0.693 (0.008) 0.896 (0.000) 0.809 (0.001) 0.817 (0.001) 0.813 (0.000) ESM-2 (150M) 0.834 (0.010) 0.758 (0.012) 0.831 (0.005) 0.823 (0.023) 0.519 (0.021) 0.636 (0.015) 0.850 (0.011) 0.586 (0.026) 0.975 (0.001) 0.948 (0.001) 0.380 (0.022) 0.540 (0.023) 0.922 (0.003) 0.714 (0.002) 0.901 (0.001) 0.820 (0.012) 0.819 (0.004) 0.819 (0.003) ESM-2 (650M) 0.850 (0.002) 0.793 (0.003) 0.847 (0.003) 0.808 (0.037) 0.606 (0.033) 0.692 (0.008) 0.857 (0.005) 0.550 (0.011) 0.964 (0.002) 0.971 (0.003) 0.088 (0.053) 0.159 (0.086) 0.928 (0.000) 0.718 (0.000) 0.900 (0.000) 0.818 (0.000) 0.820 (0.000) 0.819 (0.000) AbLang-H 0.884 (0.000) 0.850 (0.000) 0.889 (0.000) 0.927 (0.000) 0.660 (0.000) 0.771 (0.000) 0.925 (0.000) 0.737 (0.000) 0.979 (0.000) 0.937 (0.000) 0.512 (0.000) 0.662 (0.000) 0.933 (0.000) 0.756 (0.000) 0.902 (0.000) 0.821 (0.000) 0.822 (0.000) 0.821 (0.000) AbLang-L 0.811 (0.018) 0.737 (0.033) 0.824 (0.013) 0.868 (0.052) 0.448 (0.019) 0.591 (0.029) 0.914 (0.011) 0.672 (0.016) 0.978 (0.002) 0.944 (0.012) 0.482 (0.003) 0.639 (0.002) 0.933 (0.000) 0.742 (0.009) 0.900 (0.000) 0.821 (0.001) 0.811 (0.018) 0.816 (0.010) AntiBERTy 0.810 (0.004) 0.719 (0.007) 0.818 (0.004) 0.852 (0.004) 0.436 (0.035) 0.576 (0.022) 0.871 (0.003) 0.565 (0.009) 0.977 (0.002) 0.932 (0.006) 0.444 (0.010) 0.602 (0.007) 0.932 (0.002) 0.739 (0.004) 0.902 (0.001) 0.821 (0.003) 0.821 (0.003) 0.821 (0.002) AntiBERTa2 0.824 (0.001) 0.760 (0.001) 0.834 (0.000) 0.769 (0.001) 0.596 (0.000) 0.672 (0.000) 0.887 (0.006) 0.615 (0.020) 0.976 (0.001) 0.926 (0.013) 0.445 (0.032) 0.597 (0.025) 0.930 (0.001) 0.746 (0.001) 0.903 (0.002) 0.825 (0.001) 0.816 (0.020) 0.821 (0.010) AntiBERTa2-CSSP 0.845 (0.011) 0.793 (0.013) 0.850 (0.011) 0.831 (0.021) 0.591 (0.043) 0.690 (0.028) 0.916 (0.011) 0.725 (0.015) 0.981 (0.001) 0.939 (0.000) 0.548 (0.020) 0.696 (0.018) 0.939 (0.013) 0.765 (0.004) 0.906 (0.002) 0.831 (0.001) 0.822 (0.001) 0.826 (0.001) IgBert 0.817 (0.003) 0.743 (0.008) 0.817 (0.014) 0.909 (0.045) 0.399 (0.082) 0.549 (0.068) 0.821 (0.003) 0.430 (0.020) 0.969 (0.000) 0.929 (0.022) 0.247 (0.005) 0.389 (0.006) 0.921 (0.001) 0.687 (0.002) 0.896 (0.001) 0.808 (0.002) 0.820 (0.006) 0.814 (0.002) NanoBERT 0.876 (0.017) 0.829 (0.029) 0.874 (0.019) 0.873 (0.039) 0.650 (0.039) 0.745 (0.040) 0.904 (0.015) 0.687 (0.021) 0.978 (0.001) 0.948 (0.009) 0.481 (0.007) 0.638 (0.005) 0.928 (0.001) 0.719 (0.002) 0.897 (0.001) 0.812 (0.004) 0.810 (0.004) 0.811 (0.001) VHHBERT 0.800 (0.010) 0.712 (0.007) 0.811 (0.007) 0.811 (0.082) 0.454 (0.101) 0.573 (0.061) 0.873 (0.000) 0.593 (0.000) 0.976 (0.000) 0.955 (0.000) 0.426 (0.000) 0.589 (0.000) 0.921 (0.002) 0.695 (0.007) 0.890 (0.001) 0.802 (0.002) 0.794 (0.004) 0.798 (0.002)
Table 9: Performance on Thermo-seq, Thermo-tm, Affinity-seq and Affinity-score tasks. Each task reports Spearman coefficient, $R^{2}$, RMSE and MAE.
Model Thermo-seq Thermo-tm Affinity-seq Affinity-score Spearman $R^{2}$ RMSE MAE Spearman $R^{2}$ RMSE MAE Spearman $R^{2}$ RMSE MAE Spearman $R^{2}$ RMSE MAE ProtBert 0.040 (0.001) 0.003 (0.000) 20.689 (0.048) 18.767 (0.052) 0.128 (0.002) 0.013 (0.000) 24.797 (0.054) 22.886 (0.107) 0.163 (0.008) 0.044 (0.006) 4.003 (0.025) 3.358 (0.029) 0.084 (0.015) 0.030 (0.008) 1.787 (0.006) 1.296 (0.119) ESM-2 (150M) 0.389 (0.000) 0.273 (0.000) 9.946 (0.000) 7.771 (0.000) 0.301 (0.003) 0.113 (0.001) 9.255 (0.025) 7.810 (0.025) 0.170 (0.000) 0.073 (0.000) 3.935 (0.000) 3.319 (0.000) 0.063 (0.000) 0.037 (0.000) 1.842 (0.000) 1.066 (0.000) ESM-2 (650M) 0.375 (0.010) 0.154 (0.005) 10.099 (0.050) 7.683 (0.026) 0.315 (0.000) 0.096 (0.000) 9.178 (0.000) 7.842 (0.000) 0.183 (0.003) 0.082 (0.001) 3.464 (0.006) 2.779 (0.003) 0.089 (0.001) 0.052 (0.006) 1.855 (0.036) 1.267 (0.236) AbLang-H 0.566 (0.005) 0.413 (0.012) 10.482 (0.227) 8.239 (0.146) 0.533 (0.004) 0.272 (0.006) 9.916 (0.110) 7.936 (0.107) 0.151 (0.000) 0.025 (0.000) 5.031 (0.000) 4.537 (0.000) 0.106 (0.000) 0.063 (0.000) 1.898 (0.000) 1.608 (0.000) AbLang-L 0.585 (0.001) 0.412 (0.002) 9.161 (0.171) 7.173 (0.118) 0.502 (0.002) 0.264 (0.003) 8.855 (0.105) 7.143 (0.054) 0.173 (0.010) 0.053 (0.008) 4.073 (0.722) 3.316 (0.809) 0.128 (0.004) 0.058 (0.007) 2.082 (0.125) 1.821 (0.153) AntiBERTy 0.474 (0.006) 0.345 (0.009) 9.187 (0.204) 7.050 (0.119) 0.385 (0.007) 0.140 (0.003) 9.640 (0.069) 7.888 (0.063) 0.167 (0.017) 0.074 (0.017) 3.532 (0.456) 3.053 (0.544) 0.119 (0.013) 0.072 (0.003) 1.915 (0.106) 1.408 (0.194) AntiBERTa2 0.587 (0.002) 0.451 (0.000) 8.493 (0.001) 6.533 (0.000) 0.585 (0.003) 0.343 (0.002) 8.615 (0.142) 6.723 (0.145) 0.162 (0.025) 0.041 (0.007) 3.534 (0.731) 2.840 (0.844) 0.088 (0.001) 0.053 (0.009) 2.150 (0.092) 1.630 (0.409) AntiBERTa2-CSSP 0.575 (0.010) 0.459 (0.010) 8.808 (0.151) 6.708 (0.097) 0.593 (0.004) 0.361 (0.004) 8.469 (0.109) 6.839 (0.083) 0.149 (0.003) 0.067 (0.006) 3.921 (0.940) 3.307 (1.052) 0.085 (0.011) 0.066 (0.014) 2.057 (0.113) 1.418 (0.450) IgBert 0.351 (0.178) 0.156 (0.098) 10.028 (0.344) 7.873 (0.168) 0.344 (0.176) 0.144 (0.111) 9.109 (0.358) 7.544 (0.337) 0.184 (0.006) 0.082 (0.013) 3.369 (0.582) 2.670 (0.666) 0.090 (0.037) 0.065 (0.001) 1.908 (0.236) 1.285 (0.185) NanoBERT 0.440 (0.006) 0.241 (0.009) 9.590 (0.073) 7.380 (0.057) 0.298 (0.003) 0.053 (0.001) 10.011 (0.037) 7.949 (0.013) 0.061 (0.000) 0.006 (0.000) 3.518 (0.000) 2.811 (0.000) 0.118 (0.000) 0.022 (0.000) 1.895 (0.000) 1.575 (0.000) VHHBERT 0.484 (0.007) 0.390 (0.004) 9.318 (0.053) 7.540 (0.021) 0.532 (0.000) 0.310 (0.000) 8.701 (0.000) 6.883 (0.000) 0.145 (0.015) 0.044 (0.019) 3.110 (0.152) 2.248 (0.178) 0.115 (0.006) 0.070 (0.003) 2.092 (0.130) 1.667 (0.402)
Table 10: Performance on Polyreactivity prediction task. The task reports AUROC, AUPRC, Accuracy (Acc), Precision (Prec), Recall (Rec), and F1.
Model AUROC AUPRC Acc Prec Rec F1 ProtBert 0.837 (0.002) 0.844 (0.002) 0.766 (0.002) 0.787 (0.001) 0.775 (0.003) 0.781 (0.002) ESM-2 (150M) 0.833 (0.004) 0.840 (0.006) 0.764 (0.003) 0.784 (0.003) 0.775 (0.008) 0.779 (0.004) ESM-2 (650M) 0.842 (0.004) 0.847 (0.005) 0.773 (0.004) 0.783 (0.003) 0.798 (0.008) 0.791 (0.006) AbLang-H 0.831 (0.003) 0.839 (0.004) 0.765 (0.003) 0.767 (0.003) 0.805 (0.007) 0.785 (0.005) AbLang-L 0.819 (0.002) 0.823 (0.002) 0.749 (0.002) 0.771 (0.011) 0.761 (0.033) 0.765 (0.010) AntiBERTy 0.828 (0.001) 0.835 (0.001) 0.758 (0.002) 0.771 (0.013) 0.785 (0.029) 0.777 (0.008) AntiBERTa2 0.833 (0.003) 0.840 (0.004) 0.761 (0.004) 0.778 (0.005) 0.775 (0.010) 0.776 (0.007) AntiBERTa2-CSSP 0.830 (0.001) 0.838 (0.001) 0.760 (0.003) 0.779 (0.010) 0.775 (0.022) 0.776 (0.007) IgBert 0.829 (0.010) 0.836 (0.011) 0.762 (0.008) 0.784 (0.006) 0.790 (0.016) 0.786 (0.010) NanoBERT 0.815 (0.004) 0.823 (0.005) 0.747 (0.003) 0.755 (0.007) 0.785 (0.012) 0.769 (0.006) VHHBERT 0.818 (0.003) 0.823 (0.002) 0.750 (0.003) 0.767 (0.005) 0.770 (0.007) 0.768 (0.005)
Table 11: Performance on Nanobody Type prediction task. The task reports Accuracy (Acc), Precision (Prec), Recall (Rec), and F1.
Model Acc Prec Rec F1 ProtBert 0.957 (0.001) 0.995 (0.002) 0.997 (0.000) 0.996 (0.001) ESM-2 (150M) 0.994 (0.001) 0.995 (0.000) 0.988 (0.001) 0.992 (0.000) ESM-2 (650M) 0.995 (0.000) 0.996 (0.000) 0.994 (0.004) 0.995 (0.002) AbLang-H 0.997 (0.000) 0.998 (0.000) 0.997 (0.000) 0.997 (0.000) AbLang-L 0.988 (0.000) 0.990 (0.000) 0.948 (0.006) 0.966 (0.003) AntiBERTy 0.999 (0.001) 0.999 (0.001) 0.996 (0.003) 0.997 (0.001) AntiBERTa2 0.995 (0.000) 0.996 (0.000) 0.953 (0.009) 0.972 (0.005) AntiBERTa2-CSSP 0.998 (0.000) 0.998 (0.000) 0.996 (0.004) 0.997 (0.002) IgBert 0.993 (0.008) 0.994 (0.006) 0.986 (0.021) 0.990 (0.013) NanoBERT 0.988 (0.001) 0.990 (0.001) 0.947 (0.005) 0.967 (0.002) VHHBERT 0.986 (0.002) 0.989 (0.001) 0.939 (0.001) 0.963 (0.001)
NeurIPS Paper Checklist
The checklist is designed to encourage best practices for responsible machine learning research, addressing issues of reproducibility, transparency, research ethics, and societal impact. Do not remove the checklist: The papers not including the checklist will be desk rejected. The checklist should follow the references and follow the (optional) supplemental material. The checklist does NOT count towards the page limit.
Please read the checklist guidelines carefully for information on how to answer these questions. For each question in the checklist:
  • You should answer [Yes] , [No] , or [N/A] .
  • [N/A] means either that the question is Not Applicable for that particular paper or the relevant information is Not Available.
  • Please provide a short (1–2 sentence) justification right after your answer (even for NA).
The checklist answers are an integral part of your paper submission. They are visible to the reviewers, area chairs, senior area chairs, and ethics reviewers. You will be asked to also include it (after eventual revisions) with the final version of your paper, and its final version will be published with the paper.
The reviewers of your paper will be asked to use the checklist as one of the factors in their evaluation. While "[Yes] " is generally preferable to "[No] ", it is perfectly acceptable to answer "[No] " provided a proper justification is given (e.g., "error bars are not reported because it would be too computationally expensive" or "we were unable to find the license for the dataset we used"). In general, answering "[No] " or "[N/A] " is not grounds for rejection. While the questions are phrased in a binary way, we acknowledge that the true answer is often more nuanced, so please just use your best judgment and write a justification to elaborate. All supporting evidence can appear either in the main paper or the supplemental material, provided in appendix. If you answer [Yes] to a question, in the justification please point to the section(s) where related material for the question can be found.
IMPORTANT, please:
  • Delete this instruction block, but keep the section heading “NeurIPS Paper Checklist",
  • Keep the checklist subsection headings, questions/answers and guidelines below.
  • Do not modify the questions and only use the provided macros for your answers.
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The abstract and introduction clearly describe the creation of NbBench, evaluation across tasks, and standardization of benchmarks.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: Conclusion section describes key limitations, including lack of 3D structural dataset and large language model evaluation.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: The paper is empirical and does not include theoretical results.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: Task Pipeline in Results and Appendix A.1–A.4 detail the training setups, hyperparameters, evaluation metrics, and data splits.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: All datasets are public on HuggingFace, and code is publicly available on github.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification:Task Pipeline in Results and Appendix A cover task setup, model pipeline, optimizer, batch size, and splits.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: For each task and model, we conducted 3 independent runs and report the mean and standard deviation of the primary evaluation metric.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: Appendix states all experiments were conducted on a single NVIDIA A100 GPU with 80GB memory.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: All datasets have proper citations, and no human data or unethical practices are involved.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: Biological impact of each task is discussed in Benchmark Tasks (e.g., therapeutic screening, stability design).
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: No high-risk models or data (e.g., general-purpose LLMs or scraped images) are released.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: All datasets and models are cited (see Benchmark Tasks and Appendix), and public licenses respected.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: We introduce NbBench, a new benchmark suite for nanobody modeling composed of eight tasks across nine curated datasets. While the raw datasets are publicly available, we provide new standardized splits, task definitions and evaluation protocols. All documentation is included on the corrsponding github and huggingface datasets.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: The work does not involve any crowdsourcing or human participants.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: No human subjects were involved.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: No LLMs were used as core components; only specialized protein/nanobody LMs were evaluated.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.